# Proprioceptive Image: An Image Representation of Proprioceptive Data from Quadruped Robots for Contact Estimation Learning

### v1

@@ I INTRODUCTION
Deep learning has achieved remarkable success in domains where sequential or high-dimensional data can be transformed into visual representations suitable to convolutional architectures. In speech recognition, for instance, raw audio signals are often converted into spectrograms, two-dimensional time–frequency images, that serve as powerful inputs to convolutional neural networks (CNNs) and other image-based models. These approaches have yielded state-of-the-art performance across a range of tasks, from keyword spotting to end-to-end transcription [1] [2] [3]. Similarly, in cybersecurity, software files and their execution behavior have been rendered as grayscale or RGB images, where visual patterns detectable by CNNs facilitate malware detection [4]. These data-to-image transformations have unlocked new pathways by leveraging the spatial feature extraction strengths inherent to image-oriented deep learning.
Inspired by these cross-domain successes, this work investigates the transformation of proprioceptive time-series data collected from multiple proprioceptive signals of a quadruped robot into two-dimensional images that encode both temporal dynamics and robot morphology. By arranging proprioceptive signal readings according to the physical layout of the robot and incorporating time as an additional spatial dimension, the resulting images preserve inter-signal correlations and gait-dependent patterns in a form optimized for CNN processing. This approach provides a richer and more spatially coherent feature space than models trained directly on raw time-series inputs, enabling more effective learning of complex locomotion tasks. We demonstrate its potential in the context of contact estimation, where accurate prediction of foot–ground interactions is essential for stable and adaptive quadruped locomotion [5] [6].
Figure 1 illustrates three time steps of a quadruped robot walking over uneven terrain. For each time step, the corresponding proprioceptive images are shown, derived from base angular velocity from IMU, joint position from encoders, and estimated foot position proprioceptive signals. These images highlight how the representation varies across different environments, providing a compact encoding that supports learning of fundamental tasks such as foot contact estimation.
@@ I-A Related Work - 1D time series to image conversion
To apply computer vision methods to time-series analysis, Wang and Oates [7] proposed the Gramian Angular Field (GAF) and Markov Transition Field (MTF) encodings, which enable CNN-based classification. GAF constructs a Gramian matrix [8] where elements capture trigonometric relations between time intervals, offering a bijective mapping that preserves temporal correlations and allows approximate reconstruction of the original series.
In contrast, the MTF encodes the dynamic transition statistics of a time series by discretizing it into quantile bins and constructing a transition matrix over time [7]. While MTF effectively represents multi-span transition probabilities, including self-transitions along the diagonal, its performance is sensitive to the choice of bins, complicating its use with heterogeneous signal modalities. Moreover, both MTF and GAF suffer from rate invariance, producing identical images for increasing and decreasing sequences with the same rate.
The Recurrence Plot (RP) [9] captures temporal correlations by marking state recurrences, enabling the diagnosis of drift, periodicity, or chaotic dynamics. However, it is less sensitive to subtle variations, which limits its effectiveness.
Transforming time series into images such as GAF, MTF, or RP enables CNN-based spatial feature learning and has shown strong performance across benchmarks [10, 11, 12]. Multimodal fusion through meta-feature fusion, attention, or concatenation further improves accuracy by capturing complementary static, dynamic, and recurrence patterns, with successful applications in manufacturing defect detection and healthcare heartbeat classification [13, 14, 15]. Overall, fused image encodings combined with tailored CNN/attention architectures remain a practical and effective strategy for time-series classification [16, 17].
In this work, we adapt the motivation of prior studies on converting 1D time-series data into 2D image representations for feature extraction via CNNs to the context of quadruped robotics. We propose a novel encoding technique that leverages proprioceptive signal data and the robot’s morphology to generate structured image representations, enabling the robot to learn tasks such as foot contact estimation.
@@ I-B Related Work - Contact Estimation
Contact estimation for quadruped robots is a fundamental perception problem, as contact measurements are not always available [6]. However, it is a hard problem due to sensor noise, terrain variability, and other factors such as slippage.
A common method to estimate contact is to use a threshold value with the estimated or measured ground reaction forces (GRF) [18]. However, this approach often fail to generalize across robots or environments. To address these challenges, diverse approaches leverage probabilistic modeling, sensor fusion, and system dynamics. Camurri et al. [5] introduce a probabilistic method that estimates ground contact by using the GRF computed from joint positions, velocities, and efforts. Other frameworks, such as the Interacting Multiple-Model Kalman Filter (IMM-KF) [19], model the robot’s movement as a switched system where each mode represents a different set of feet in contact. Alternatively, Maravgakis et al. [20] propose a probabilistic technique that uses Kernel Density Estimation (KDE) over batches of IMU data to compute the probability of stable six degree-of-freedom contact. Other methods include the Dual $\beta$-Kalman Filter [21], which confronts the challenges of foot slippage and variable leg length.
Traditional model-based and thresholding approaches to contact estimation remain sensitive to noise and require empirical tuning, limiting robustness on lightweight robots or varying terrains. To overcome these challenges, recent work has moved toward data-driven techniques. For instance, Youm et al. [22] use a Neural Measurement Network to output contact probability, enhancing IEKF performance under slip conditions. Similarly, Lee et al. [23] estimate contact probabilistically from dynamics and joint torques, with a neural compensator correcting residual errors. Beyond these hybrid strategies, fully learning-based methods have also emerged. For instance, Lin et al. [24] proposed Contact-CNN, a multi-modal deep model that relies solely on proprioceptive signals (joint encoders, IMU, kinematics) to classify foot–ground contact across diverse gaits and terrains, avoiding explicit contact sensing altogether.
Ordonez-Apraez et al. [25] introduced Discrete Morphological Symmetries (DMSs), a group-theoretic symmetry groups of dynamical systems. Building on this, they proposed ECNN, a variant of Contact-CNN that applies sparse equivariant convolutions and batched augmentation to enforce hard equivariance, making large architectures tractable and improving leg-level F1 and test loss. Also, MI-HGNN [26] introduces a morphology-informed heterogeneous graph neural network that constructs a graph from a robot’s kinematic structure (nodes being the base/joints/feet, and edges representing links) to constrain learning for contact perception problems. Empirically, MI-HGNN attains substantially higher contact-detection accuracy (mean 87.7% vs 78.8% from ECNN [25]) on the Contact Data Set [24].
@@ I-C Contributions
  • We propose a CNN architecture that is able to learn PI information and perform contact estimation for quadruped robots.
@@ II METHODOLOGY
In this section, we present the four key components of the PI formulation and explain how they are fused to construct the final PI representation, which is tailored to the robot’s morphology. Each component captures essential temporal and statistical characteristics of a single proprioceptive state and encodes them into a sub-image through well-defined mathematical transformations. Specifically, it is extracted and encoded into image form the following elements from the signal data stream: the slope dynamics in Section II-B, spike patterns in Section II-C, modified GAF polar distance in Section II-D, and the global-aware local shift in Section II-E. Figure 2 shows how to generate the Proprioceptive Image from proprioceptive signal data.
@@ II-A Preliminaries
Input Values: Let $s=\{s_{i}\}^{T}_{i=1}$ be a discrete 1D time series sampled from a proprioceptive signal (estimated or measured by a sensor). At each timestamp $t$, we define a sliding window with length size $w$ as $x_{t}=[s_{t-w+1},\dots,s_{t}]\in\mathbb{R}^{w}$.
@@
Global Bounds: The global bounds are defined as the maximum and minimum values of the proprioceptive signals outputs $\mathcal{G}=[s^{global}_{min},s^{global}_{max}]$ (usually obtained from the sensor datasheet). $\mu_{\mathcal{G}}$ is defined as the average of the global bounds $\mu_{\mathcal{G}}=(s^{global}_{max}+s^{global}_{min})/2$. The normalized sliding window sequence, computed with respect to the global bounds, is defined as $x_{t}^{\mathcal{G}}=[s_{t-w+1}^{\mathcal{G}},\dots,s_{t}^{\mathcal{G}}]\in\mathbb{R}^{w}$, where each $s_{i}^{\mathcal{G}}\in[-1,1]$ represents the globally normalized signal value at time $i$ computed as:
@@
Local Range: For each proprioceptive signal of interest, it is computed a pair of normalized deviation measures $\delta$. These values quantify how far the current local operating range $\mathcal{Y}=[s^{local}_{min},s^{local}_{max}]$ of the signal is shifted relative to its global reference range. The local range is estimated from a sliding window by taking the lower and upper percentiles (10th and 90th) and expanding them by a small margin to ensure robustness to noise. Given the global average $\mu_{\mathcal{G}}$ and global bounds $\mathcal{G}$, the normalized deviation $\delta=[\delta_{min},\delta_{max}]$ is defined as:
@@ II-B Slope Dynamics
Given the temporal window of scalar values $x_{t}$, we encode three dynamical descriptors — mean slope, jerk, and peak density — into a spatial Gaussian pattern modulated by concentric ripples. More precisely, we fit a linear model $x_{i}=\beta_{t}i+b$ over the window $w$ to compute the slope $\beta_{t}\in\mathbb{R}$ using least-squares and clipped to be [1,1][-1,1]:
@@ II-D Polar Distance
In this sub-image, we employ a modified version of the GAF representation. While the standard GAF inherently captures static temporal information through its unique polar encoding, it suffers from ambiguity in distinguishing between increasing and decreasing trends. To overcome this limitation, we replace the values along the main diagonal of the GAF matrix with the globally normalized sliding window sequence $x_{t}^{\mathcal{G}}$, subsequently scaled to the pixel intensity range [0, 255]. This substitution embeds directional information directly into the matrix diagonal values, enhancing the interpretability of the encoded temporal dynamics. Figure 5 shows three representative proprioceptive images corresponding to different input conditions. Eq. 15 describes the GAF function:
$D=GAF(x_{t},x_{t}^{\mathcal{G}})\in\mathbb{R}^{w\times w}$ (15)
@@ II-F Contact Estimation
Following [24], contact estimation is framed as a multi-class classification problem, where each binary foot–ground configuration corresponds to a discrete class. To address this, we propose ConcatCNN, a multi-image convolutional architecture that jointly encodes and fuses proprioceptive image representations. Each input image is processed by a shared CNN encoder with two convolutional layers (kernel size $3\times 3$, padding 1), each followed by Group Normalization [28] and Leaky ReLU activation [29], with a $2\times 2$ max-pooling operation for downsampling. Leaky ReLU mitigates the risk of dead neurons in sparse proprioceptive inputs, while ReLU activations are retained in the fully connected layers for stability and efficiency. Group Normalization was chosen over Batch Normalization for its robustness to small batch sizes and heterogeneous signals streams, ensuring consistent feature scaling. The shared encoder outputs feature maps of size $16\times 5\times 5$ per image, which are concatenated into a fused representation of size $16N\times 5\times 5$, where $N$ denotes the number of proprioceptive images (PIs) provided as input. This representation is refined by a convolutional fusion layer (kernel size $3\times 3$, 32 channels), flattened, and passed through a two-layer fully connected network (hidden dimension 128, ReLU activation) with dropout [30] for regularization. The final output layer predicts the contact state across the predefined classes. Overall, ConcatCNN enables efficient joint feature extraction and fusion by leveraging parameter sharing in the encoder and channel-wise concatenation for cross-signal integration, thereby enhancing the robustness of contact state classification. The complete architecture is depicted in Fig. 7, which highlights the shared encoders, fusion stage, and final prediction layers.
Figure 7: Architecture of ConcatCNN for quadruped contact estimation. Each proprioceptive image is encoded by a shared CNN, and the resulting features are stacked and fused to jointly learn from multiple proprioceptive signals. In the case of fusing trunk and leg PIs, the concatenated four legs PI are separated. The fused representation is flattened and mapped through fully connected layers to predict discrete foot–ground contact states.
@@ III-A Real-World Dataset Results
We also evaluated our approach using the same metrics as [25] [26], namely the foot-wise binary F1-score, the averaged F1-score (mean of the four leg-specific F1-scores), and the contact state accuracy over 16 possible states. The latter is a stricter metric than averaged accuracy, as a prediction is counted as correct only when the contact state of all four feet is simultaneously correctly predicted. To compute this metric for our model, the foot-wise predictions were converted into a 16-state representation. In addition, the real-world dataset provided by [24] was transformed into PIs using a sliding window of size $w=10$, consistent with the results reported in Tab. I. Unlike the original setup, however, our model was trained on the combined training and validation sets and evaluated on the combined test sets. The experimental results, presented in Fig. 8, demonstrate that our method outperforms all state-of-the-art data-driven methods (MI-HGNN [26], ECNN [25], CNN-aug [25], and CNN [24]) in the Legs-Avg F1 (from 0.9356 of MI-HGNN to 0.9732) and overall model accuracy (from 87.7% of MI-HGNN to 94.5%) with smaller standard deviation. These results show the potential of this representation by obtaining better accuracy with a less complex NN architecture with relation to other methods [25] [26].

### v2

@@ I INTRODUCTION
Deep learning has been highly successful when sequential or high-dimensional signals are converted into visual representations that suit convolutional architectures. In speech recognition, raw audio is commonly transformed into spectrograms, i.e., 2D time–frequency images, enabling CNN-based models to achieve state-of-the-art results, from keyword spotting to end-to-end transcription [5, 1, 22]. Similarly, in cybersecurity, software files and their execution behavior have been mapped to grayscale or RGB images, where CNNs exploit visual patterns for malware detection [13]. These data-to-image mappings leverage the spatial feature learning strengths of image-based deep networks.
Motivated by these successes, we transform quadruped proprioceptive time-series signals into 2D images that encode both temporal dynamics and robot morphology. Signals are arranged according to the robot’s physical layout, with time incorporated as an additional spatial dimension, preserving inter-signal correlations and gait-dependent structure in a CNN-friendly format. This yields a richer and more spatially coherent feature space than learning directly from raw time series, and we demonstrate its effectiveness on foot contact estimation, a key component for stable and adaptive locomotion [4, 18].
Figure 1 shows three time steps of a quadruped traversing uneven terrain, along with the corresponding proprioceptive images derived from IMU base angular velocity, encoder joint positions, and estimated foot positions. The resulting patterns vary across environments while providing a compact representation that supports learning tasks such as contact estimation.
@@ I-A Related Work - 1D time series to image conversion
To enable CNN-based classification on time-series data, Wang and Oates [26] proposed the Gramian Angular Field (GAF) and Markov Transition Field (MTF) encodings. GAF constructs a Gramian matrix [9] whose entries capture trigonometric relations between time indices, yielding a bijective mapping that preserves temporal correlations and supports approximate reconstruction.
In contrast, MTF represents temporal dynamics through transition statistics: the signal is discretized into quantile bins and a time-indexed transition matrix is formed [26]. Although MTF captures multi-span transition probabilities (including self-transitions along the diagonal), it is sensitive to the binning choice, which complicates its use with heterogeneous modalities. Moreover, both MTF and GAF exhibit rate invariance, producing identical images for increasing and decreasing sequences with the same rate.
Recurrence Plots (RP) [6] encode temporal structure by marking state recurrences, supporting the diagnosis of drift, periodicity, and chaotic dynamics, but they are less responsive to subtle variations.
Overall, converting 1D time series into 2D encodings (GAF, MTF, RP) enables spatial feature learning with CNNs and has shown strong performance across benchmarks [33, 20, 21]. Performance can be further improved via multimodal fusion (meta-features, attention, or concatenation), which captures complementary static, dynamic, and recurrence cues, with applications ranging from manufacturing defect detection to heartbeat classification [30, 29, 27]. As such, fused image encodings paired with tailored CNN/attention architectures remain a practical and effective strategy for time-series classification [25, 11].
In this work, we bring this paradigm to quadruped robotics by converting proprioceptive time-series signals into structured 2D representations informed by robot morphology, enabling CNN-based feature extraction for tasks such as foot contact estimation.
@@ I-B Related Work - Contact Estimation
Contact estimation for quadruped robots is a fundamental perception problem, since contact measurements are not always available [18]. Yet it remains challenging due to sensor noise, terrain variability, and effects such as slippage.
A common baseline estimates contact by thresholding measured or estimated ground reaction forces (GRF) [7], but it often fails to generalize across robots and environments. To improve robustness, several works combine probabilistic modeling, sensor fusion, and system dynamics. Camurri et al. [4] estimate contact probabilistically using GRF computed from joint positions, velocities, and efforts. Menner and Berntorp [17], instead, cast locomotion as a switched system, where each mode corresponds to a different set of feet in contact. Maravgakis et al. [16] use Kernel Density Estimation (KDE) over IMU batches to infer the probability of stable 6-DoF contact, while the Dual $\beta$-Kalman Filter [32] explicitly addresses slippage and variable leg length.
Despite these advances, model-based and thresholding approaches remain noise-sensitive and typically require empirical tuning, limiting robustness on lightweight robots and changing terrains. Consequently, recent work has shifted toward data-driven techniques. Youm et al. [31] learn a contact-probability measurement model via a Neural Measurement Network to improve IEKF performance under slip, and Lee et al. [10] combine probabilistic torque/dynamics-based estimation with a neural compensator for residual errors. Beyond these hybrid strategies, fully learning-based methods have also emerged. For example, Lin et al. [12] propose Contact-CNN, a multi-modal network that relies only on proprioceptive signals (joint encoders, IMU, kinematics) to classify foot–ground contact across diverse gaits and terrains, removing the need for explicit contact sensing.
Ordonez-Apraez et al. [19] introduced Discrete Morphological Symmetries (DMSs), a group-theoretic symmetry groups of dynamical systems. Building on this, they proposed ECNN, a variant of Contact-CNN that applies sparse equivariant convolutions and batched augmentation to enforce hard equivariance, making large architectures tractable and improving leg-level F1 and test loss. Also, MI-HGNN [3] introduces a morphology-informed heterogeneous graph neural network (GNN) that constructs a graph from a robot’s kinematic structure (nodes being the base/joints/feet, and edges representing links) to constrain learning for contact perception problems. Empirically, MI-HGNN attains substantially higher contact-detection accuracy (mean 87.7% vs 78.8% from ECNN [19]) on the Contact Data Set [12].
@@ I-C Contributions
  • We propose a CNN architecture that is able to learn from PI information and perform contact estimation for quadruped robots.
@@ II METHODOLOGY
In this section, we present the four key components of the PI formulation and explain how they are fused to construct the final PI representation, which is tailored to the robot’s morphology. Each component captures essential temporal and statistical characteristics of a single proprioceptive state and encodes them into a sub-image through well-defined mathematical transformations. Specifically, it is extracted and encoded into image form using the following elements from the signal data stream: the slope dynamics in Section II-B, spike patterns in Section II-C, modified GAF polar distance in Section II-D, and the global-aware local shift in Section II-E. Figure 2 shows how to generate the Proprioceptive Image from proprioceptive signal data.
@@ II-A Preliminaries
Input Values: Let $s=\{s_{i}\}^{T}_{i=1}$ be a discrete 1D time series sampled from a proprioceptive signal (estimated or measured by a sensor). At each timestamp $t$, we define a sliding window of length $w$ as $x_{t}=[s_{t-w+1},\dots,s_{t}]\in\mathbb{R}^{w}$. The goal is to convert each key component into a $w\times w$ image and concatenate them to form a $2w\times 2w$ image.
@@
Global Bounds: The PI’s representation uses as reference the global bounds of the input signal. The global bounds are defined as the maximum and minimum values of the proprioceptive signals outputs $\mathcal{G}=[s^{global}_{min},s^{global}_{max}]$ (usually obtained from the sensor datasheet). $\mu_{\mathcal{G}}$ is defined as the average of the global bounds $\mu_{\mathcal{G}}=(s^{global}_{max}+s^{global}_{min})/2$. The normalized sliding window sequence, computed with respect to the global bounds, is defined as $x_{t}^{\mathcal{G}}=[s_{t-w+1}^{\mathcal{G}},\dots,s_{t}^{\mathcal{G}}]\in\mathbb{R}^{w}$, where each $s_{i}^{\mathcal{G}}\in[-1,1]$ represents the globally normalized signal value at time $i$ computed as:
@@
Local Range: For each proprioceptive signal of interest, we compute a pair of normalized deviation measures $\delta$. These values quantify how far the current local operating range $\mathcal{Y}=[s^{local}_{min},s^{local}_{max}]$ of the signal is shifted relative to its global reference range. The local range is estimated from a sliding window by taking the lower and upper percentiles (10th and 90th) and expanding them by a small margin to ensure robustness to noise. Given the global average $\mu_{\mathcal{G}}$ and global bounds $\mathcal{G}$, the normalized deviation $\delta=[\delta_{min},\delta_{max}]$ is defined as:
@@ II-B Slope Dynamics
Given the temporal window of scalar values $x_{t}$, we encode three dynamical descriptors — mean slope, jerk, and peak density — into a spatial Gaussian blob modulated by concentric ripples. More precisely, we fit a linear model $x_{i}=\beta_{t}i+b$ over the window $w$ to compute the slope $\beta_{t}\in\mathbb{R}$ using least-squares and clipped to be [1,1][-1,1]:
@@ II-D Polar Distance
In this sub-image, we employ a modified version of the GAF representation. GAF maps the input signal to an angular coordinate $\phi_{t}=arccos(x^{\mathcal{G}}_{t})$ and its matrix entries are defined as $GAF=\cos(\phi_{i}+\phi_{j}),(i,j)=1,\dots,w$. While the standard GAF inherently captures static temporal information through its unique polar encoding, it suffers from ambiguity in distinguishing between increasing and decreasing trends. To overcome this limitation, we replace the values along the main diagonal of the GAF matrix with the globally normalized sliding window sequence $x_{t}^{\mathcal{G}}$, subsequently scaled to the pixel intensity range [0, 255]. This substitution embeds directional information directly into the matrix diagonal values, enhancing the interpretability of the encoded temporal dynamics. Figure 5 shows three representative proprioceptive images corresponding to different input conditions. Eq. 15 describes the GAF function:
$\centering D=\begin{bmatrix}x^{\mathcal{G}}_{1}&\cos(\phi_{1}+\phi_{2})&\dots&\cos(\phi_{1}+\phi_{w})\\ \cos(\phi_{2}+\phi_{1})&x^{\mathcal{G}}_{2}&\dots&\vdots\\ \vdots&\vdots&\ddots&\vdots\\ \cos(\phi_{w}+\phi_{1})&\dots&\dots&x_{w}^{\mathcal{G}}\end{bmatrix}\@add@centering$ (15)
@@ II-F Contact Estimation
Following [12], contact estimation is framed as a multi-class classification problem, where each binary foot–ground configuration corresponds to a discrete class. To address this, we propose ConcatCNN, a multi-image convolutional architecture that jointly encodes and fuses proprioceptive image representations. The input modalities are joint positions (rad), joint velocities (rad/s), base linear acceleration ($m/s^{2}$), base angular velocity (rad/s), foot positions (m), and foot velocities (m/s), all expressed in the base frame when applicable. Each input image is processed by a shared CNN encoder with two convolutional layers (kernel size $3\times 3$, padding 1), each followed by Group Normalization (GN) [28] and LReLU activation [15], with a $2\times 2$ max-pooling operation for downsampling. LReLU mitigates the risk of dead neurons in sparse proprioceptive inputs, while ReLU activations are retained in the fully connected layers for stability and efficiency. GN was chosen over Batch Normalization for its robustness to small batch sizes and heterogeneous signals streams, ensuring consistent feature scaling. The shared encoder outputs feature maps of size $16\times 5\times 5$ per image, which are concatenated into a fused representation of size $16N\times 5\times 5$, where $N$ denotes the number of PIs provided as input. This representation is refined by a convolutional fusion layer (kernel size $3\times 3$, 32 channels), flattened, and passed through a two-layer fully connected network (hidden dimension 128, ReLU activation) with dropout [8] for regularization. The final output layer predicts the contact state across the predefined classes. Overall, ConcatCNN enables efficient joint feature extraction and fusion by leveraging parameter sharing in the encoder and channel-wise concatenation for cross-signal integration, thereby enhancing the robustness of contact state classification. The complete architecture is depicted in Fig. 7, which highlights the shared encoders, fusion stage, and final prediction layers.
Figure 7: Architecture of ConcatCNN for quadruped contact estimation. Each proprioceptive image is encoded by a shared CNN, and the resulting features are stacked and fused to jointly learn from multiple proprioceptive signals. In the case of fusing trunk and leg PIs, the concatenated four legs PI are separated. The fused representation is flattened and mapped through fully connected layers to predict discrete foot–ground contact states as a classification label.
@@ III-A Real-World Dataset Results
We also evaluated our approach using the same metrics as [19] [3], namely the foot-wise binary F1-score, the averaged F1-score (mean of the four leg-specific F1-scores), and the contact state accuracy over 16 possible states. The latter is a stricter metric than averaged accuracy, as a prediction is counted as correct only when the contact state of all four feet is simultaneously correctly predicted. To compute this metric for our model, the foot-wise predictions were converted into a 16-state representation. In addition, the real-world dataset provided by [12] was transformed into PIs using a sliding window of size $w=10$, consistent with the results reported in Tab. I. Unlike the original setup, however, our model was trained on the merged training and validation splits (treated as a single training set) and evaluated on the merged test splits. The experimental results, presented in Fig. 8, demonstrate that our method outperforms all state-of-the-art data-driven methods (MI-HGNN [3], ECNN [19], CNN-aug [19], and CNN [12]) in the Legs-Avg F1 (from 0.9356 of MI-HGNN to 0.9732) and overall model accuracy (from 87.7% of MI-HGNN to 94.5%) with smaller standard deviation. These results show the potential of this representation to achieve higher accuracy with a less complex NN architecture compared to other methods [19][3].